Please help.
A person scores 81 on a test of verbal ability and 6.4 on a test of quantitative ability. For the verbal ability test, the mean for people in general is 50 and the standard deviation is 20. For the quantitative ability test, the mean for people in general is 0 and the standard deviation is 5. Which is the person’s stronger ability: verbal or quantitative? Explain your answer to a person who has never had a course in statistics.

Answer by stanbon(75887)

A person scores 81 on a test of verbal ability and 6.4 on a test of quantitative ability. For the verbal ability test, the mean for people in general is 50 and the standard deviation is 20. For the quantitative ability test, the mean for people in general is 0 and the standard deviation is 5. Which is the person’s stronger ability: verbal or quantitative? Explain your answer to a person who has never had a course in statistics.81) =
-----------------------------------------
Standardize both scores by converting them to z-scores
z(81) = (81-50)/2 = 31/20 = 1.55
----------
z(6.4) = (6.4-0)/5 = 1.28
--------------------------------------
Conclusion: The verbal ability test is the higher.
--------------------------------------
Cheers,
Stan